# Acoustic Analogy of Quantum Baldin Sum Rule for Optimal Causal Scattering

### v1

@@ Abstract
journal: a peer-review journal
@@
Introduction.— The transmission loss (TL) of a soundproofing wall with thickness $L$ increases by 6 dB for each doubling of either the effective mass density per unit area ($\rho_{\mathrm{eff}}L$) or the angular frequency $\omega$, known as the ’mass law’: $\mathrm{TL}\approx 20\log_{10}\left(\frac{\omega\rho_{\text{eff}}L}{2Z_{0}}\right)$, where the characteristic impedance of the background fluid is given by $Z_{0}=\sqrt{\rho_{0}K_{0}}$ (with $\rho_{0}$ and $K_{0}$ representing the mass density and bulk modulus of the background fluid, respectively). In 2000, Liu et al. liu2000locally challenged this principle by introducing lead spheres encased in a soft rubber and epoxy matrix, which exhibit deep-subwavelength dipole resonances characterized by dynamically negative mass density. This negative property lee2016origin can lead to a bandgap with strong backscattering, even when the scatterer unit is significantly smaller than the excitation wave’s wavelength, thereby violating the mass law. Fang et al. fang2006ultrasonic further identified the monopole counterpart of negative bulk modulus through an engineered Helmholtz resonator array. In essence, metamaterials can concentrate modes within a specific band of interest, to create large range of dynamic properties, by leveraging the density of states from other frequencies monticone2013cloaked, wang2025seven, landi2018acoustic. This enables exciting applications such as wave focusing oh2023engineering, cloaking xu2021transformation, imaging zhu2011holey, tweezing xu2024acoustofluidic, and perfect absorption qu2022microwave, yang2025acoustic. However, the mass law disregards the casual dispersion of materials and could be inapplicable at excessively low frequencies (stiffness control region) or excessively high frequencies (where higher-order modes emerge). The well-known Rozanov limit addresses the causality-induced dispersion constraint on one-port absorption bandwidth rozanov2000ultimate, acher2009fundamental, yang2017sound, yet it assumes a transmission-forbidden boundary. Thus, an open research question remains on the universal casual bound governing transmission loss: how the allocation of local-band spectral modes dictate the global wave transport?
@@
To address this problem, we draw inspiration from micro-scale causal property of the quantum theory, the Baldin sum rule BALDIN1960310, specifically a fundamental result in quantum field theory that connects the electric ($\alpha_{e}$) and magnetic ($\alpha_{m}$) polarizabilities of nucleons to the integral of the photo-absorption extinction cross-section $\sigma_{\mathrm{ext}}$: $\int_{\nu_{0}}^{\infty}\sigma_{\mathrm{ext}}{d\nu/\nu^{2}}=2\pi^{2}(\alpha_{e}+\alpha_{m})$ (natural unit), where $\nu_{0}$ denotes the threshold energy to trigger nuclear light-matter interaction. As shown in Fig. (1a), by measuring the far-field forward scattering amplitude $f(\theta=0)$, optical theorem gustafsson2012optical, a result of scattering unitarity, gives $\sigma_{\mathrm{ext}}=(4\pi/k_{0})\mathrm{Im}[f(\theta=0)]$ ($k_{0}$ is the incident wavenumber). So, nucleon structure parameters can be directly extracted from scattering-based experimental data schumacher2005polarizability, holstein2014hadron.
@@
Model definitions.— Acoustic metamaterials are commonly realized as periodic subwavelength structures or duct-embedded resonators for effective control of wave propagation. Accordingly, we focus on a one-dimensional (1D) scattering model [Fig. (1b)], distinct from the three-dimensional (3D) spherical scattering scenario [Fig. (1a)]. We define the complex transmission coefficient $T(\omega)$ as the ratio of the transmitted total field $p_{t}$ to the incident field $p_{i}$ across a structure of thickness $L$ (where $L\ll\lambda$), whose effective density and longitudinal modulus denote $\rho_{\mathrm{eff}}$ and $M_{\mathrm{eff}}$, respectively (see a homogenization method in Ref. yang2014homogenization). Thus, the normalized forward scattering amplitude is given by $T(\omega)-1$, derived from the scattered field $p_{s}=p_{t}-p_{i}$, while $R(\omega)$ denotes the reflection coefficient or backward scattering amplitude. The extinction cross-section area is defined as the sum of the absorption term $(1-|R(\omega)|^{2}-|T(\omega)|^{2})S_{0}$] and the scattered backward and forward cross-sections $(|R(\omega)|^{2}+|T(\omega)-1|^{2})S_{0}$], where $S_{0}$ denotes the periodic wavefront/port area or duct cross-section area. Therefore, we reduce the extinction cross-section $\sigma_{\mathrm{ext}}$ (normalized by $S_{0}$) to
@@
which can be regarded as a 1D optical theorem. Since $T(\omega)-1$ is analytic in the upper half-plane of frequency (refer to Ref. labelSM, Sec. S1), we apply the Kramers–Kronig (KK) relations waters2000applicability to obtain ($\mathcal{P}$ denotes the principal value)
@@
Meanwhile, the analytical expression of $T(\omega)$ is labelSM
@@
Analytical validation.— We investigated typical cases of ideal point scatterers, with closed-form dispersion models of $\rho_{\mathrm{eff}}(\omega)$ and $M_{\mathrm{eff}}(\omega)$. Mathematically, we have analytically obtained perfect agreement between $\int_{0}^{\infty}\sigma_{\mathrm{ext}}(\omega)\frac{d\omega}{\omega^{2}}$ and $\Gamma$ for all cases (see derivations in Ref. labelSM, Sec. S1).
Numerical validation.— Ideal analytical models neglect higher-order modes. Here, the numerical examples we investigated include both a monopole Helmholtz resonator fang2006ultrasonic and a lead-core dipole resonator liu2000locally that incorporate multiple elastic modes [see Fig. (2)]. Using material parameters from the original studies, we performed finite element simulations (FEM) to calculate $\sigma_{\mathrm{ext}}(\omega)$, with water set as the background media (see Ref. labelSM, Sec. S2 for simulation details). As shown in Fig. (2c), the monopole resonator—mounted on a duct—exhibits two subwavelength peaks: the first corresponds to the Helmholtz resonance ($L/\lambda\approx 1/5$), and the second to the half-wavelength Fabry–Pérot resonance ($L/\lambda\approx 1/2$) within the defined structural region of length $L$. For $L/\lambda>1$, oscillations arise due to higher-order Fabry–Pérot resonances and associated phase modulation. Similarly, the dipole resonator array, shows multiple peaks ($L/\lambda$ as small as $1/200$), attributed to the elastic modal behavior of composite materials still2013soft. After extracting $M_{\mathrm{eff}}(\omega)$ and $\rho_{\mathrm{eff}}(\omega)$ spectra via $R(\omega)$ and $T(\omega)$ using the method in Ref. groby2021analytical, we observe that the corresponding lineshapes near resonance follow the Lorentz model, as shown in the insets of Fig. (2c).
@@
which approaches the bound $\Gamma$ as $\omega\to\infty$ $\gamma(\omega)/\Gamma\to 1$ for both cases in Fig. (2d)]. We adopted the effective medium theory of composite materials zhou2009analytic, milton2022theory to calculate $\Gamma$ with the static normalized effective density ${\rho_{\mathrm{eff}}(0)}/{\rho_{0}}$ and compressibility ${K_{0}}/{M_{\mathrm{eff}}(0)}$ [see Ref. labelSM, Sec. S2]. By comparing Figs. (2c) and (2d), we can clearly see the local-band contribution in $\gamma(\omega)$ from the low-frequency local resonances to the higher-order Fabry–Pérot resonances.
Design strategy.— The acoustic Baldin sum rule serves as a predictive tool: by maximally suppressing the low-frequency $\sigma_{\mathrm{ext}}$, scattering resources are redistributed to higher frequencies, broadening the operational bandwidth [Fig. (1c)]. To quantify this spectral shaping effect, we combine Eq. (6) and Eq. (2) and then compute the derivative of $\gamma(\omega)$ (${\sigma_{\mathrm{ext}}(\omega)}/{\omega^{2}}$) at the low frequency limit
$\frac{d\gamma(\omega)}{d\omega}=\frac{\omega^{2}L^{2}}{2c_{0}^{2}}\left[\left(\frac{\rho_{\mathrm{eff}}(0)}{\rho_{0}}\right)^{2}+\left(\frac{K_{0}}{K_{\mathrm{eff}}(0)}\right)^{2}\right]+\mathcal{O}(\omega^{2}),$ (8)
where we replace $M_{\mathrm{eff}}(\omega)$ by $K_{\mathrm{eff}}(\omega)$ (bulk modulus) for the fluid-based scatterers without shear modes. Because $\left(\frac{\rho_{\mathrm{eff}}(0)}{\rho_{0}}\right)^{2}+\left(\frac{K_{0}}{K_{\mathrm{eff}}(0)}\right)^{2}\geq 2{\frac{\rho_{\mathrm{eff}}(0)}{\rho_{0}}}{\frac{K_{0}}{K_{\mathrm{eff}}(0)}}$, the coefficient of $\omega^{2}$ can be optimally minimized as $L^{2}/c_{\mathrm{eff}}^{2}(0)$, if the static effective and compressibility can be matched
@@
We implement this principle using a Fano resonator, the simplest system coupling a discrete monopole resonance (featured by $K_{0}/K_{\text{eff}}(\omega)=\frac{\alpha(2c_{0}/L)}{\omega_{m}^{2}-\omega^{2}-i\beta\omega}$ ) to a continuum dipole background ($\frac{\rho_{\mathrm{eff}}(\omega)}{\rho_{0}}\approx\frac{2c_{0}}{L}\sqrt{b}$). The different combination of $\rho_{\mathrm{eff}}(\omega)$ and $K_{\mathrm{eff}}(\omega)$ can be interpreted as the interference of two oscillators. By taking $k_{\mathrm{eff}}L\ll 1$ in Eq. (4), we can obtain the transmission coefficient for Fano resonance
@@
In fact, Ref. goffaux2002evidence derived an exactly same formula of Eq. (10) via a different approach. According to Eq. (9), to minimize the low-frequency $\omega^{2}$ coefficient of $\sigma_{\mathrm{ext}}$, it yields $\sqrt{b}=\alpha/\omega_{m}^{2}$ (optimal Fano scattering condition), which results in a maximum degree of asymmetry in the transmission lineshape [refer to Ref. labelSM, Sec. S4].
Experimental verification.— To validate the sum rule and its design principle, we fabricated three acoustic resonators within a two-port air duct: a foam liner, a Helmholtz resonator, and a Fano resonator [Figs. (3a–c)]. A useful principle qu2025duality to determine their effective properties is
@@
where $\varphi$ is the perforation ratio of the narrowed channel (with $\Delta L$ accounting for the near-field end correction qu2024analytical), and $V_{\mathrm{eff}}$ is the total effective volume of the scatterer, defined with Wood’s formula milton2022theory, ge2025causal. In Table (A2) of End Matter, we present our simplified model of the three designed resonators with the following features
@@
    Foam liner: $K_{0}/K_{\mathrm{eff}}(0)=11.65$ (due to the volume of additional porous material installed on the sidewalls of the pipe), and $\rho_{\mathrm{eff}}(0)/\rho_{0}\approx 1$ (fully ventilated with $\varphi\approx 1$). The foam liner can be treated as a lossy monopole resonator. Porous material region was modelled with Johnson-Champoux-Allard model yang2017sound.
@@
    Helmholtz resonator: $K_{0}/K_{\mathrm{eff}}(0)=8.32$ (with the same air volume as the foam liner, but lower effective compressibility due to the dominant adiabatic process at low frequencies ge2025causal), and $\rho_{\mathrm{eff}}(0)/\rho_{0}\approx 1$. The Helmholtz resonator can be approximated as a lossless monopole resonator.
@@
By using simulation and experiments (details in Ref. labelSM, Sec. S3 and S4), in Fig. (3d), we confirmed superior sound insulation performance of Fano resonator with optimal scattering condition [Eq. (9)], with an average measured $\langle\mathrm{TL}\rangle=21.3\,\mathrm{dB}$ in the frequency range of $1098\,\mathrm{Hz}$$6174\,\mathrm{Hz}$. However, for Helmholtz resonator, $\langle\mathrm{TL}\rangle=18.6\,\mathrm{dB}$ in the range $960\,\mathrm{Hz}$$2332\,\mathrm{Hz}$, and for the foam liner, $\langle\mathrm{TL}\rangle=13.8\,\mathrm{dB}$ in the range $960\,\mathrm{Hz}$$3156\,\mathrm{Hz}$. The Fano resonator’s total $V_{\mathrm{eff}}$ is less than that of other two cases, but the TL performance is better due to the design guided by sum rule.
Figure 3: The experimental validation via airborne sound resonators in ducts. (a) Foam liner (monopole type). (b) Helmholtz resonator (monopole type). (c) Fano resonator (coupled monopole-dipole type). The size of the oscillator sphere represents the amount of dissipation, and the length of the line represents the resonant frequency. Dashed lines represent 2D rotational symmetry axes, and solid lines represent hard boundaries. (d) The simulated (solid lines) and measured (circles) transmission spectra. (e) The cumulative distribution function $\gamma(\omega)$. The inset shows the calculated $K_{0}/K_{\mathrm{eff}}(0)+\rho_{\mathrm{eff}}(0)/\rho_{0}$.
In Fig. (3e), we confirmed that the optimally interfered Fano resonator squeezes more $\sigma_{\mathrm{ext}}$ resources to a higher frequency band of interest, as characterized by the lowest $\gamma(\omega)$ curve, resulting in the exceptional TL bandwidth shown in Fig. (3d). For all cases, with broadband enough spectra data, the measured ultimate bound verified the correctness of acoustic Baldin sum rule: $\gamma(2\pi\times 6500\mathrm{Hz})/\Gamma$ approaches unity, as shown in Fig. (3e). We further showed that the measured $T(\omega)$ with both phase and amplitude, shown in Fig. (4) of End Matter, follows the Kramers-Kronig relations as well.
Equation (6) is only an asymptotic form for low frequencies. To establish the family of causal scattering functions over the broad band, we introduce Padé approximant model friedland2012control with order $m\times n$]:
@@
where $|T(\omega)|\leq 1$ with $m\leq n$ (passivity) and causality prohibits the presence of poles in the upper half-plane $(\operatorname{Im}(\omega)\geq 0)$, under the constraint of Routh–Hurwitz criterion friedland2012control. If $\omega\to 0$ (Taylor series), $2[1-T(\omega)]=\left(-ia_{1}\right)\omega+\left(a_{1}b_{1}-a_{2}\right)\omega^{2}+i\mathcal{O}(\omega^{3})$, so, $a_{1}=2\Gamma/\pi$ (fixed by sum rule’s bound). Except $a_{1}$, other coefficients (real-valued) contribute to the full degree of freedom of spectral shaping. We fitted the three cases in Fig. (3) with a Padé model of minimal order ($m=3$, $n=3$) to capture the low-frequency feature of Fano resonance as per Eq. (10) [see the fitting details in Ref. labelSM, Sec. S4]. As shown in Fig. (3e), $\gamma(\omega)$ from Padé models (dashed lines) also validated the bound of sum rule and spectral shaping effect.
Discussion.— Usually, to achieve ultra-broadband wave manipulation, multiple resonators must be integrated, thus introducing undesired algorithmic complexity and fabrication challenges. Here, counter-intuitively, our Fano structure is remarkably simple, representing the minimal physical elements required to achieve causally optimal scattering. A sum-rule-based Figure of Merit further confirms that this design ranks among the highest for broadband ventilated silencers wang2023meta, nguyen2020broadband, jia2025acoustic, lu2025ultra, mei2025reconfigurable, dong2021ultrabroadband, xu2024broadband, fu2025ultrabroadband, jimenez2017rainbow [see Fig. (5) in End Matter]. Although many theoretical versions of sum rules have been proposed for classical acoustic meng2022fundamental, norris2018integral and electromagnetic padilla2024fundamental, gustafsson2010sum waves, we reveal the missing link to the Baldin sum rule in quantum field theory, and achieve its direct experimental verification and first application in acoustics. Although we are only discussing the 1D plane wave scatterings, the relevant theoretical framework can be easily extended to 2D and 3D systems ramachandran2023bandwidth, martin2024acoustic.
Beyond acoustics, this work bridges a conceptual gap, revealing the classical counterpart of a fundamental quantum scattering constraint. The underlying framework, described by the family of causal Padé approximants described by Eq. (12), is universal. It provides a practical design strategy applicable not only to suppressing waves (as in absorbers qu2024analytical, qu2025duality, yang2025acoustic, super-scatterers ramachandran2023bandwidth) but to enhancing wave transport (e.g., cloaking monticone2013cloaked, impedance matching layers dong2025soft, dong2020bioinspired), and even to spectrum customization applications wang2023meta, qu2023reverberation.

### v2

@@ Abstract
journal: Physical Review Letters
@@
Introduction.— The transmission loss (TL) of a soundproofing wall with thickness $L$ increases by 6 dB for each doubling of either the effective mass density per unit area ($\rho_{\mathrm{eff}}L$) or the angular frequency $\omega$, known as the ’mass law’: $\mathrm{TL}\approx 20\log_{10}\left(\frac{\omega\rho_{\text{eff}}L}{2Z_{0}}\right)$, where the characteristic impedance of the background fluid is given by $Z_{0}=\sqrt{\rho_{0}K_{0}}$ (with $\rho_{0}$ and $K_{0}$ representing the mass density and bulk modulus of the background fluid, respectively). In 2000, Liu et al. Liu et al. [2000] challenged this principle by introducing lead spheres encased in a soft rubber and epoxy matrix, which exhibit deep-subwavelength dipole resonances characterized by dynamically negative mass density. This negative property Lee and Wright [2016] can lead to a bandgap with strong backscattering, even when the scatterer unit is significantly smaller than the excitation wave’s wavelength, thereby violating the mass law. Fang et al. Fang et al. [2006] further identified the monopole counterpart of negative bulk modulus through an engineered Helmholtz resonator array. In essence, metamaterials can concentrate modes within a specific band of interest, to create large range of dynamic properties, by leveraging the density of states from other frequencies Monticone and Alu [2013], Wang et al. [2025], Landi et al. [2018]. This enables exciting applications such as wave focusing Oh et al. [2023], cloaking Xu and Chen [2021], imaging Zhu et al. [2011], tweezing Xu et al. [2024a], and perfect absorption Qu and Sheng [2022], Yang et al. [2025]. However, the mass law disregards the casual dispersion of materials and could be inapplicable at excessively low frequencies (stiffness control region) or excessively high frequencies (where higher-order modes emerge). The well-known Rozanov limit addresses the causality-induced dispersion constraint on one-port absorption bandwidth Rozanov [2000], Acher et al. [2009], Yang and Sheng [2017], yet it assumes a transmission-forbidden boundary. Thus, an open research question remains on the universal casual bound governing transmission loss: how the allocation of local-band spectral modes dictate the global wave transport?
@@
To address this problem, we draw inspiration from micro-scale causal property of the quantum theory, the Baldin sum rule Baldin [1960], specifically a fundamental result in quantum field theory that connects the electric ($\alpha_{e}$) and magnetic ($\alpha_{m}$) polarizabilities of nucleons to the integral of the photo-absorption extinction cross-section $\sigma_{\mathrm{ext}}$: $\int_{\nu_{0}}^{\infty}\sigma_{\mathrm{ext}}{d\nu/\nu^{2}}=2\pi^{2}(\alpha_{e}+\alpha_{m})$ (natural unit), where $\nu_{0}$ denotes the threshold energy to trigger nuclear light-matter interaction. As shown in Fig. (1a), by measuring the far-field forward scattering amplitude $f(\theta=0)$, optical theorem Gustafsson et al. [2012], a result of scattering unitarity, gives $\sigma_{\mathrm{ext}}=(4\pi/k_{0})\mathrm{Im}[f(\theta=0)]$ ($k_{0}$ is the incident wavenumber). So, nucleon structure parameters can be directly extracted from scattering-based experimental data Schumacher [2005], Holstein and Scherer [2014].
@@
Model definitions.— Acoustic metamaterials are commonly realized as periodic subwavelength structures or duct-embedded resonators for effective control of wave propagation. Accordingly, we focus on a one-dimensional (1D) scattering model [Fig. (1b)], distinct from the three-dimensional (3D) spherical scattering scenario [Fig. (1a)]. We define the complex transmission coefficient $T(\omega)$ as the ratio of the transmitted total field $p_{t}$ to the incident field $p_{i}$ across a structure of thickness $L$ (where $L\ll\lambda$), whose effective density and longitudinal modulus denote $\rho_{\mathrm{eff}}$ and $M_{\mathrm{eff}}$, respectively (see a homogenization method in Ref. Yang et al. [2014]). Thus, the normalized forward scattering amplitude is given by $T(\omega)-1$, derived from the scattered field $p_{s}=p_{t}-p_{i}$, while $R(\omega)$ denotes the reflection coefficient or backward scattering amplitude. The extinction cross-section area is defined as the sum of the absorption term $(1-|R(\omega)|^{2}-|T(\omega)|^{2})S_{0}$] and the scattered backward and forward cross-sections $(|R(\omega)|^{2}+|T(\omega)-1|^{2})S_{0}$], where $S_{0}$ denotes the periodic wavefront/port area or duct cross-section area. Therefore, we reduce the extinction cross-section $\sigma_{\mathrm{ext}}$ (normalized by $S_{0}$) to
@@
which can be regarded as a 1D optical theorem (derivations based on energy conservation are available in Ref. 1, Sec. S1). Since $T(\omega)-1$ is analytic in the upper half-plane of frequency, we apply the Kramers–Kronig (KK) relations Waters et al. [2000] to obtain ($\mathcal{P}$ denotes the principal value)
@@
Meanwhile, the analytical expression of $T(\omega)$ is 1
@@
Analytical validation.— We investigated typical cases of ideal point scatterers, with closed-form dispersion models of $\rho_{\mathrm{eff}}(\omega)$ and $M_{\mathrm{eff}}(\omega)$. Mathematically, we have analytically obtained perfect agreement between $\int_{0}^{\infty}\sigma_{\mathrm{ext}}(\omega)\frac{d\omega}{\omega^{2}}$ and $\Gamma$ for all cases (see derivations in Ref. 1, Sec. S1).
Numerical validation.Ideal analytical models, as given in Eq. (4), neglect higher-order modes (in waveguide scattering) or diffraction modes (in planar periodic samples). However, the sum rule does not preclude the existence of dynamic modes beyond the effective medium description, since all scattering states remain orthogonal.
Here, the numerical examples we investigated include both a monopole Helmholtz resonator Fang et al. [2006] and a lead-core dipole resonator Liu et al. [2000] that incorporate higher-order fluid-based or elastic modes [see Fig. (2)]. Using material parameters from the original studies, we performed finite element simulations (FEM) to calculate $\sigma_{\mathrm{ext}}(\omega)$, with water set as the background media (see Ref. 1, Sec. S2 for simulation details). As shown in Fig. (2c), the monopole resonator—mounted on a duct—exhibits two subwavelength peaks: the first corresponds to the Helmholtz resonance ($L/\lambda\approx 1/5$), and the second to the half-wavelength Fabry–Pérot resonance ($L/\lambda\approx 1/2$) within the defined structural region of length $L$. For $L/\lambda>1$, oscillations arise due to higher-order Fabry–Pérot resonances and associated phase modulation. Similarly, the dipole resonator array, shows multiple peaks ($L/\lambda$ as small as $1/200$), attributed to the elastic modal behavior of composite materials Still et al. [2013]. After extracting $M_{\mathrm{eff}}(\omega)$ and $\rho_{\mathrm{eff}}(\omega)$ spectra via $R(\omega)$ and $T(\omega)$ using the method in Ref. Groby et al. [2021], we observe that the corresponding lineshapes near resonance follow the Lorentz model, as shown in the insets of Fig. (2c).
@@
which approaches the bound $\Gamma$ as $\omega\to\infty$ $\gamma(\omega)/\Gamma\to 1$ for both cases in Fig. (2d)]. We adopted the effective medium theory of composite materials Zhou and Hu [2009], Milton [2022] to calculate $\Gamma$ with the static normalized effective density ${\rho_{\mathrm{eff}}(0)}/{\rho_{0}}$ and compressibility ${K_{0}}/{M_{\mathrm{eff}}(0)}$ [see Ref. 1, Sec. S2]. By comparing Figs. (2c) and (2d), we can clearly see the local-band contribution in $\gamma(\omega)$ from the low-frequency local resonances to the higher-order Fabry–Pérot resonances.
Design strategy.— The acoustic Baldin sum rule serves as a predictive tool: by maximally suppressing the low-frequency $\sigma_{\mathrm{ext}}$, scattering resources are redistributed to higher frequencies, broadening the operational bandwidth [Fig. (1c)]. To quantify this spectral shaping effect, we combine Eq. (6) and Eq. (2) and then compute $d\gamma(\omega)/d\omega$ (${\sigma_{\mathrm{ext}}(\omega)}/{\omega^{2}}$) at the low frequency limit
$\frac{d\gamma(\omega)}{d\omega}=\frac{L^{2}}{2c_{0}^{2}}\left[\left(\frac{\rho_{\mathrm{eff}}(0)}{\rho_{0}}\right)^{2}+\left(\frac{K_{0}}{K_{\mathrm{eff}}(0)}\right)^{2}\right]+\mathcal{O}(\omega^{2}),$ (8)
where we replace $M_{\mathrm{eff}}(\omega)$ by $K_{\mathrm{eff}}(\omega)$ (bulk modulus) for the fluid-based scatterers without shear modes. Because $\left(\frac{\rho_{\mathrm{eff}}(0)}{\rho_{0}}\right)^{2}+\left(\frac{K_{0}}{K_{\mathrm{eff}}(0)}\right)^{2}\geq 2{\frac{\rho_{\mathrm{eff}}(0)}{\rho_{0}}}{\frac{K_{0}}{K_{\mathrm{eff}}(0)}}$, the coefficient of $\omega^{2}$ in $\sigma_{\text{ext}}(\omega)$ can be optimally minimized as $L^{2}/c_{\mathrm{eff}}^{2}(0)$, if the static effective and compressibility can be matched
@@
This is a necessary but not sufficient condition for optimal scattering. An additional consideration is to ensure that $\frac{\rho_{\mathrm{eff}}(0)}{\rho_{0}}=\frac{K_{0}}{K_{\mathrm{eff}}(0)}\neq 1$ ($c_{\mathrm{eff}}\neq c_{0}$), so that the static properties contrast with those of the background fluid as a precondition for dynamic modal control. To make this strategy clear, we construct a Fano resonator, the simplest system coupling a discrete monopole resonance (featured by $K_{0}/K_{\text{eff}}(\omega)=\frac{\alpha(2c_{0}/L)}{\omega_{m}^{2}-\omega^{2}-i\beta\omega}$ ) to a continuum dipole background ($\frac{\rho_{\mathrm{eff}}(\omega)}{\rho_{0}}\approx\frac{2c_{0}}{L}\sqrt{b}$). The different combination of $\rho_{\mathrm{eff}}(\omega)$ and $K_{\mathrm{eff}}(\omega)$ can be interpreted as the interference of two oscillators. By taking $k_{\mathrm{eff}}L\ll 1$ in Eq. (4), we can obtain the transmission coefficient for Fano resonance
@@
In fact, Ref. Goffaux et al. [2002] derived an exactly same formula of Eq. (10) via a different approach. According to Eq. (9), to minimize the low-frequency $\omega^{2}$ coefficient of $\sigma_{\mathrm{ext}}$, it yields $\sqrt{b}=\alpha/\omega_{m}^{2}$ (optimal Fano scattering condition), which results in a maximum degree of asymmetry in the transmission lineshape [refer to Ref. 1, Sec. S4].
Experimental verification.— To validate the sum rule and its design principle, we fabricated three acoustic resonators within a two-port air duct: a foam liner, a Helmholtz resonator, and a Fano resonator [Figs. (3a–c)]. A useful principle Qu et al. [2025] to determine their effective properties is
@@
where $\varphi$ is the perforation ratio of the narrowed channel (with $\Delta L$ accounting for the near-field end correction Qu et al. [2024]), and $V_{\mathrm{eff}}$ is the total effective volume of the scatterer, defined with Wood’s formula Milton [2022], Ge et al. [2025]. In Table (A2) of End Matter, we present our simplified model of the three designed resonators with the following features
@@
    Foam liner: $K_{0}/K_{\mathrm{eff}}(0)=11.65$ (due to the volume of additional porous material installed on the sidewalls of the pipe), and $\rho_{\mathrm{eff}}(0)/\rho_{0}\approx 1$ (fully ventilated with $\varphi\approx 1$). The foam liner can be treated as a lossy monopole resonator. Porous material region was modelled with Johnson-Champoux-Allard model Yang and Sheng [2017].
@@
    Helmholtz resonator: $K_{0}/K_{\mathrm{eff}}(0)=8.32$ (with the same air volume as the foam liner, but lower effective compressibility due to the dominant adiabatic process at low frequencies Ge et al. [2025]), and $\rho_{\mathrm{eff}}(0)/\rho_{0}\approx 1$. The Helmholtz resonator can be approximated as a lossless monopole resonator.
@@
By using simulation and experiments (details in Ref. 1, Sec. S3 and S4), in Fig. (3d), we confirmed superior sound insulation performance of Fano resonator with optimal scattering condition [Eq. (9)], with an average measured $\langle\mathrm{TL}\rangle=21.3\,\mathrm{dB}$ in the frequency range of $1098\,\mathrm{Hz}$$6174\,\mathrm{Hz}$. However, for Helmholtz resonator, $\langle\mathrm{TL}\rangle=18.6\,\mathrm{dB}$ in the range $960\,\mathrm{Hz}$$2332\,\mathrm{Hz}$, and for the foam liner, $\langle\mathrm{TL}\rangle=13.8\,\mathrm{dB}$ in the range $960\,\mathrm{Hz}$$3156\,\mathrm{Hz}$. The Fano resonator’s total $V_{\mathrm{eff}}$ is less than that of other two cases, but the TL performance is better due to the design guided by sum rule.
Figure 3: The experimental validation via airborne sound resonators in ducts. (a) Foam liner (monopole type). (b) Helmholtz resonator (monopole type). (c) Fano resonator (coupled monopole-dipole type). The size of the oscillator sphere represents the amount of dissipation, and the length of the line represents the resonant frequency. Dashed lines represent 2D rotational symmetry axes, and solid lines represent hard boundaries. (d) The simulated (solid lines) and measured (circles) transmission spectra. (e) The cumulative distribution function $\gamma(\omega)$. The inset shows the calculated values of $K_{0}/K_{\mathrm{eff}}(0)$ and $\rho_{\mathrm{eff}}(0)/\rho_{0}$, along with their ratios and the simulation-extracted values. The experimental $\gamma(\omega)$ below 100 Hz is complemented by low-frequency asymptotic forms derived from simulations (see Ref. 1 Sec. S4 for details).
In Fig. (3e), we confirmed that the optimally interfered Fano resonator squeezes more $\sigma_{\mathrm{ext}}$ resources to a higher frequency band of interest, as characterized by the lowest $\gamma(\omega)$ curve, resulting in the exceptional TL bandwidth shown in Fig. (3d). We emphasize that the sum rule is a universal constraint on the relative bandwidth, because Eq. (1) can alternatively be written in a dimensionless form: $\int_{0}^{\infty}\sigma_{\mathrm{ext}}(\lambda)\,d\left(\frac{\lambda}{L}\right)=\pi^{2}\left(\frac{K_{0}}{M_{\mathrm{eff}}(0)}+\frac{\rho_{\mathrm{eff}}(0)}{\rho_{0}}\right)$. If $|T|\to 0$ (high TL) over the operating wavelength range $\lambda\in[\lambda_{1},\lambda_{2}]$, then $\int_{\lambda_{1}/L}^{\lambda_{2}/L}\sigma_{\mathrm{ext}}(\lambda)\,d\left(\frac{\lambda}{L}\right)\approx 2\frac{\lambda_{2}-\lambda_{1}}{L}$, which reveals how $\sigma_{\mathrm{ext}}$ is related to the TL bandwidth. For all cases, with broadband enough spectra data, the measured ultimate bound verified the correctness of acoustic Baldin sum rule: $\gamma(2\pi\times 6500\mathrm{Hz})/\Gamma$ approaches unity, as shown in Fig. (3e). We further showed that the measured $T(\omega)$ with both phase and amplitude, shown in Fig. (4) of End Matter, follows the Kramers-Kronig relations as well.
Equation (6) is only an asymptotic form for low frequencies. To overcome the implicit nature of the KK relations and establish an explicit family of causal scattering functions over a broad band, we introduce a Padé approximant model Friedland [2012] of order $m\times n]$:
@@
where $|T(\omega)|\leq 1$ with $m\leq n$ (passivity) and causality prohibits the presence of poles in the upper half-plane $(\operatorname{Im}(\omega)\geq 0)$, under the constraint of Routh–Hurwitz criterion Friedland [2012]. If $\omega\to 0$ (Taylor series), $2[1-T(\omega)]=\left(-ia_{1}\right)\omega+\left(a_{1}b_{1}-a_{2}\right)\omega^{2}+i\mathcal{O}(\omega^{3})$, so, $a_{1}=2\Gamma/\pi$ (fixed by sum rule’s bound). Except $a_{1}$, other coefficients (real-valued) contribute to the full degree of freedom of spectral shaping. We fitted the three cases in Fig. (3) with a Padé model of minimal order ($m=3$, $n=3$) to capture the low-frequency feature of Fano resonance as per Eq. (10) [see the fitting details in Ref. 1, Sec. S4]. As shown in Fig. (3e), $\gamma(\omega)$ from Padé models (dashed lines) also validated the spectral shaping effect.
Discussion.— Usually, to achieve ultra-broadband wave manipulation, multiple resonators must be integrated, thus introducing undesired algorithmic complexity and fabrication challenges. Here, counter-intuitively, our Fano structure is remarkably simple, representing the minimal physical elements required to achieve causally optimal scattering. A sum-rule-based Figure of Merit further confirms that this design ranks among the highest for broadband ventilated silencers Wang et al. [2023], Nguyen et al. [2020], Jia et al. [2025], Lu et al. [2025], Mei et al. [2025], Dong et al. [2021], Xu et al. [2024b], Fu and Hang [2025], Jiménez et al. [2017] [see Fig. (5) in End Matter]. Although many theoretical versions of sum rules have been proposed for classical acoustic Meng et al. [2022], Norris [2018] and electromagnetic Padilla et al. [2024], Gustafsson and Sjöberg [2010] waves, we reveal the missing link to the Kramer-Kronig relations and Baldin sum rule in quantum field theory, and achieve its direct experimental verification and first application in acoustics. Although we are only discussing the 1D plane wave scatterings, the relevant theoretical framework can be easily extended to 2D and 3D systems Ramachandran and Rajagopal [2023], Martin [2024].
Beyond acoustics, this work bridges a conceptual gap, revealing the classical counterpart of a fundamental quantum scattering constraint. The underlying framework, described by the family of causal Padé approximants described by Eq. (12), is universal. It provides a practical design strategy applicable not only to suppressing waves (as in absorbers Qu et al. [2024, 2025], Yang et al. [2025], super-scatterers Ramachandran and Rajagopal [2023]) but to enhancing wave transport (e.g., cloaking Monticone and Alu [2013], impedance matching layers Dong et al. [2025, 2020]), and even to spectrum customization applications Wang et al. [2023], Qu et al. [2023].